OpenStudy (anonymous):

If Susie types at a rate of 24 words every 40 seconds, how many words will she type in 20 minutes? 720 48 12 2,000

OpenStudy (studygurl14):

Again, direct proportion. w = words s = seconds k = constant (aka rate) \(\large w=ks\) "a rate of 24 words every 40 seconds" Plug in 24 for w, and 40 for s, and solve for k

OpenStudy (anonymous):

divide first plug in then divide again right?

OpenStudy (studygurl14):

not exactly...you need to divide to find k, then plug in values, then...you will need to multiply to find the answer. You'll see. What did you get for k?

OpenStudy (anonymous):

oh sorry

OpenStudy (studygurl14):

It's fine @judtpleasegahd . What value did you get for k?

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

there is a simple method for this using unitary method it is given that she writes 24 words in 40 sec => 24/40 words in 1 second -------------- (1) now we know that 20 minutes = 1200 seconds using (1) 1200 seconds = 24 / 40 *1200 = 0.6 *1200 = 720 so she can write 720 words in 20 minutes

OpenStudy (studygurl14):

correct. now we plug in 0.6 for k. Here's the tricky part. It says in 20 minutes. But we've been dealing with seconds. So we need to convert 20 minuted into seconds so that we can plug it into the equation for s and solve for w

OpenStudy (studygurl14):

@ashsek you are correct, but please do not write the answer to a question when someone is currently explaining how to do the problem. It is against the code of conduct.

OpenStudy (anonymous):

multiply

OpenStudy (studygurl14):

"Don't just provide the answer to a problem when someone else is in the middle of helping!" From the Code of Conduct http://openstudy.com/code-of-conduct @ashsek

OpenStudy (studygurl14):

@judtpleasegahd look at when @ashsek put. That is the correct answer, and at least he/she provided steps. :)

OpenStudy (anonymous):

ok ty

OpenStudy (studygurl14):

:)

OpenStudy (anonymous):

sorry @StudyGurl14 i didnt knew anything about all this code of conduct as i am new here

OpenStudy (studygurl14):

It's ok @ashsek This tutorial can help you get started. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8